# Vision Transformers Don’t Need Trained Registers

### v2

@@ 1 Introduction
In this work, we argue that while registers are indeed useful, the models don’t need to be retrained with them. Instead, we show that registers can be added post hoc, without any additional training. To do this, we first investigate the mechanism underlying the emergence of high-norm tokens. We identify a sparse set of neurons – register neurons – that create the outlier tokens by contributing high-norm values to them.
By directly editing the activation maps of the register neurons during the forward pass of the network, we can induce the formation of the high-norm activations at arbitrary token positions (see Figure 1). We use it to create new register tokens at test-time, even in models that were never trained with them, by appending new tokens and activating the register neurons in their positions.
@@ 2 Related work
High-norm tokens in transformers. Transformers tend to create high-norm tokens during their internal computation, when trained on language tasks (Xiao et al., 2024) or on vision tasks (Darcet et al., 2024). For language models, Xiao et al. (2024) showed that transformers allocate excessive attention to these high-norm tokens, and named them “attention sinks.” Sun et al. (2024) demonstrated that “attention sinks” emerge due to previous massive neuron activations. Yona et al. (2025) linked the emergence of “attention sinks” to the inability of language models to repeatedly generate a single token, and suggested a test-time fix by zeroing out the relevant activated neurons. For vision models, Darcet et al. (2024) demonstrated the emergence of high-norm tokens in low-informative areas, and suggested retraining the model with extra tokens (“registers”) for removing these high-norm artifacts from the image tokens. Wang et al. (2024) used a simpler fine-tuning approach in DINOv2 to avoid complete re-training. We aim to shift outliers at test-time without any additional training by editing the activations of the neurons that are responsible for their emergence.
@@ 3.1 How do outliers emerge in ViTs?
Highly activated neurons activate on all outlier locations. We present the full activation maps of three neurons with high activations in Figure 4. These neurons sparsely activate on all outlier positions, not just the maximum outlier position. This suggests that they are not position-specific but, rather, are responsible for outliers generally. Given these observations, we develop a method to automatically detect these neurons in the next section.
@@ 3.2 Register neurons
Detecting register neurons. Based on our hypothesis, we formulate a simple algorithm to find register neurons in Algorithm 1. Our method finds neurons whose average activation at outlier positions is consistently high across images. To detect outlier positions (FindOutliers), we follow Darcet et al. (2024) and output the positions for which the norms of the corresponding image tokens are above a predefined threshold. Our algorithm searches for neurons in layers before outliers start, by setting the $top\_layer$ parameter. We output $top\_k$ neurons with the highest average activations. For OpenCLIP, we set $top\_layer=5$, the outlier threshold at 75, and $top\_k=10$. Since register neurons manage where outliers emerge, we can also intervene upon them to move outliers to arbitrary positions.
@@ 4 Adding registers at test-time
Moving outliers to added tokens. As outlier patches lose their local patch information (Darcet et al., 2024), it is undesirable to have them within the image since it may affect downstream performance. Previous work has suggested retraining ViTs with extra tokens to remove high-norm artifacts and attention sinks. However, retraining existing ViTs is expensive, so we propose to add an extra input token and move the outliers there with register neurons.
@@
Test-time registers hold global information. We verify that test-time registers encapsulate global image information (e.g., image class) similarly to learned registers (Darcet et al., 2024). To assess this, we perform linear probing on both learned and patched registers for classification on ImageNet (Deng et al., 2009), CIFAR-10, and CIFAR-100 (Krizhevsky et al., 2009). We also compare their performance to the CLS token and a token that corresponds to a central patch in the image. As shown in Table 1, the classification accuracy of test-time registers closely matches that of trained registers and is slightly lower than the CLS token performance. These results suggest that test-time registers, like their learned counterparts, effectively capture global image-level information.
@@ 5.3 Unsupervised object discovery
Test-time registers improve unsupervised object discovery. We report correct localization (corloc) scores in Table 5. We find that LOST performance improves significantly over the base model, on features computed with our method for DINOv2 (increase of ${\sim}{21}$ in corloc). Adding the test-time register closes the gap between the baseline model and a model with trained registers, reaching within ${\sim}{0{-}2}$ corloc. This suggests that the test-time register mimics the role of trained registers on this task. However, we note that test-time registers only marginally affect the results for OpenCLIP, a phenomenon also observed in Darcet et al. (2024) with trained registers. Further analysis can be found in Section A.8.
@@ 5.5 Typographic attacks
We demonstrate here additional evidence of our ability to shift high-norm artifacts arising from register neurons. We strategically place high-norm artifacts within an image to mask out adversarial patches while preserving semantic content. This intervention is highly localized and relies on activating only a small fraction of neurons to produce targeted changes in the image representation.
Azuma & Matsui (2023) showed that CLIP is vulnerable to typographic attacks, where images can be misclassified based on written text in the image rather than the depicted object (Figure 8).
@@
Results. Qualitative results in Figure 8 show that shifting the activation maps of register neurons early in the model’s computation causes outlier patches to later appear corresponding to the text area. In Table 7, we present the attack success rate and find that it significantly drops after our intervention. This indicates that our method masks a portion of the model’s internal representation of the image without harming the semantic content. Moreover, our results suggest that patch information at outliers is, indeed, lost from the model computation as opposed to being transferred to other image patches. We also present results of an alternative masking procedure that operates in the input space. We set all of the pixels of patches containing text content to the mean pixel value of that region. Our intervention matches this alternative method’s performance with only a sparse intervention. Whereas ${\sim}10\%$ of the input space must be masked in order to have a causal effect on CLIP’s output, our intervention repurposes CLIP’s existing internal mechanisms, modifying roughly $0.02\%$ of all neurons. The effectiveness of this sparse, targeted modification in activation space highlights the influence of register neurons in guiding CLIP’s visual representation and output behavior.
@@ 6 Discussion, limitations, and future work
While our analysis shows that we can steer the location of high-norm tokens, we only addressed one component type that is responsible for their creation – neurons, while neglecting other possible elements that can contribute to their formation (e.g., attention layers or input tokens). The edited tokens result in slight performance differences with their learned counterpart (Table 1), suggesting that the test-time registers are not fully equivalent to the high-norm tokens. Finally, similarly to Darcet et al. (2024), we mostly focus on individual models (CLIP-ViT-B/16 and DINOv2-ViT-L/14). We do not present results on other, less commonly used, pretrained ViTs, and leave it for future work.
@@ Appendix A Appendix
We investigate the impact of the number of test-time registers and the initialization strategy in Sections A.1 and A.2. Additional qualitative attention visualizations are provided in Section A.3, followed by further analysis of OpenCLIP register neuron activations in Sections A.4 and A.5. We apply the analysis from Section 3 to DINOv2 in Section A.6. Additional results from the experiments in Section 5 are included in Sections A.7 and A.8. Finally, we provide additional VLM results with test-time registers in Section A.9.
@@ A.3 Attention maps after intervening upon register neurons
A.4 Full experiment results on test-time registers in OpenCLIP
@@
A.5 Additional experiments on register neurons in OpenCLIP
@@
A.6 Investigating outliers in DINOv2
@@
A.7 Extended Dense Prediction Results
@@
A.8 Additional LOST Object Discovery Results

### v4

@@ 1 Introduction
In this work, we argue that while registers are indeed useful, the models don’t need to be retrained with them. Instead, we show that registers can be added post hoc, without any additional training. To do this, we first investigate the mechanism underlying the emergence of high-norm tokens. We identify a sparse set of neurons – register neurons – that create the outlier tokens by contributing high-norm values to them. By directly editing the activation maps of the register neurons during the forward pass of the network, we can induce the formation of the high-norm activations at arbitrary token positions (see Figure 1). We use it to create new register tokens at test-time, even in models that were never trained with them, by appending new tokens and activating the register neurons in their positions.
@@ 2 Related work
High-norm tokens in transformers. Transformers tend to create high-norm tokens during their internal computation, when trained on language tasks (Xiao et al., 2024) or on vision tasks (Darcet et al., 2024). For language models, Xiao et al. (2024) showed that transformers allocate excessive attention to these high-norm tokens, and named them “attention sinks.” Sun et al. (2024) demonstrated that “attention sinks” emerge due to previous massive neuron activations. Yona et al. (2025) linked the emergence of “attention sinks” to the inability of language models to repeatedly generate a single token, and suggested a test-time fix by zeroing out the relevant activated neurons. For vision models, Darcet et al. (2024) demonstrated the emergence of high-norm tokens in low-informative areas and suggested retraining the model with extra tokens (“registers”) to remove these high-norm tokens from the image patches. Wang et al. (2024) used a simpler fine-tuning approach in DINOv2 to avoid complete re-training. More similar to our training-free approach, Nakamura et al. (2024) suppressed artifacts during inference by modifying last-layer attention, showing that it improved image clustering performance. In contrast, we remove high-norm outliers at their source by editing the activations of the neurons that create them, aiming to eliminate their effect on later computations.
@@ 3.1 How do outliers emerge in ViTs?
Highly activated neurons activate on all outlier locations. We present the full activation maps of three neurons with high activations on the top outlier in Figure 4, with more shown in Section A.4. These neurons sparsely activate on all outlier positions, not just the maximum outlier position. This suggests that they are not position-specific but, rather, are responsible for outliers generally. Given these observations, we develop a method to automatically detect these neurons in the next section.
@@ 3.2 Register neurons
Detecting register neurons. Based on our hypothesis, we formulate a simple algorithm to find register neurons in Algorithm 1. Our method finds neurons whose average activation at outlier positions is consistently high across images. To detect outlier positions (FindOutliers), we follow Darcet et al. (2024) and output the positions for which the norms of the corresponding image tokens are above a predefined threshold. Our algorithm searches for neurons in layers before outliers start, set by the $top\_layer$ parameter. We output $top\_k$ neurons with the highest average activations. For OpenCLIP, we set $top\_layer=5$, the outlier threshold at 75, and $top\_k=10$. Since register neurons set where outliers emerge, we can also intervene upon them to move outliers to arbitrary positions.
@@ 4 Adding registers at test-time
Moving outliers to added tokens. As outlier patches lose their local patch information (Darcet et al., 2024), it is undesirable to have them within the image since it may affect downstream performance. Previous work has suggested retraining ViTs with extra tokens to remove high-norm artifacts and attention sinks. However, retraining existing ViTs is expensive, so we propose to add an extra input token and move the outliers there with register neurons. Importantly, these outliers are essential to the model’s internal computations and must appear either within a token or as an attention bias (Section A.11). For instance, zeroing out the register neurons’ activation maps instead of moving the outliers causes OpenCLIP ViT-B/16’s zero-shot ImageNet performance to drop $\sim$20%.
@@
Test-time registers hold global information. We verify that test-time registers encapsulate global image information (e.g., image class) similarly to trained registers (Darcet et al., 2024). To assess this, we perform linear probing on both trained and test-time registers for classification on ImageNet (Deng et al., 2009), CIFAR-10, and CIFAR-100 (Krizhevsky et al., 2009). We also compare their performance to the CLS token and a token that corresponds to a central patch in the image. As shown in Table 1, the classification accuracy of test-time registers closely matches that of trained registers and is slightly lower than the CLS token performance. These results suggest that test-time registers, like their learned counterparts, effectively capture global image-level information.
@@ 5.3 Unsupervised object discovery
Test-time registers improve unsupervised object discovery. We report correct localization (corloc) scores in Table 5. We find that LOST performance improves significantly over the base model on features computed with our method for DINOv2 (increase of ${\sim}{21}$ corloc). Adding the test-time register closes the gap between the baseline model and a model with trained registers, reaching within ${\sim}{0{-}2}$ corloc. This suggests that the test-time register mimics the role of trained registers on this task. However, we note that test-time registers only marginally affect the results for OpenCLIP, a phenomenon also observed in Darcet et al. (2024) with trained registers. Further analysis can be found in Section A.9.
@@ 5.5 Typographic attacks
We demonstrate here additional evidence of our ability to shift high-norm artifacts arising from register neurons. Azuma & Matsui (2023) showed that CLIP is vulnerable to typographic attacks, where images can be misclassified based on written text in the image rather than the depicted object (Figure 8). We strategically place high-norm artifacts within an image to mask out adversarial patches while preserving semantic content. This intervention is highly localized and relies on activating only a small fraction of neurons to produce targeted changes in the image representation.
@@
Results. Qualitative results in Figure 8 show that shifting the activation maps of register neurons early in the model’s computation causes outlier patches to later appear corresponding to the text area. Table 7 presents the attack success rate and find that it significantly drops after our intervention, indicating that our method masks part of the model’s internal representation without harming semantic content. Our results suggest that patch information at outliers is lost from the model computation rather than transferred elsewhere. We also present results of an alternative masking procedure in the input space, setting areas corresponding to text to the mean pixel value of the region. Our intervention matches this method’s performance using only a sparse modification – repurposing CLIP’s internal mechanisms by modifying roughly $0.02\%$ of all neurons, compared to masking ${\sim}10\%$ of the input. This highlights the influence of register neurons in guiding CLIP’s visual representation and output behavior.
@@ 6 Discussion, limitations, and future work
While our analysis shows that we can steer the location of high-norm tokens, we only addressed one component type that is responsible for their creation – neurons, while neglecting other possible elements that can contribute to their formation, such as attention layers or positional encodings (Yang et al., 2024). The edited tokens result in slight performance differences with their learned counterpart (Table 1), suggesting that the test-time registers are not fully equivalent to the high-norm tokens. Finally, similarly to Darcet et al. (2024), we mostly focus on individual models (CLIP-ViT-B/16 and DINOv2-ViT-L/14). We do not present results on other, less commonly used, pretrained ViTs, and leave it for future work.
@@ Appendix A Appendix
We investigate the impact of the number of test-time registers and the initialization strategy in Sections A.1 and A.2. Additional qualitative attention visualizations are provided in Section A.3, followed by further analysis of OpenCLIP register neuron activations in Sections A.5 and A.6. We apply the analysis from Section 3 to DINOv2 in Section A.7. Additional results from the experiments in Section 5 are included in Sections A.8 and A.9. Finally, we provide additional VLM results with test-time registers in Section A.10.
@@ A.3 Attention maps after intervening upon register neurons
A.4 Additional activation maps of register neurons
While our primary focus is on layer 6 of OpenCLIP (Section 3.1), we also include additional activation maps highlighting neurons that strongly activate at the top outlier location. As shown in Figure 14, the most active neurons from earlier layers produce activation maps that align closely with the output patch norms. This consistent alignment supports our hypothesis in Section 3.2 that a small number of sparsely activating neurons determine outlier locations even before they fully emerge. In OpenCLIP, we observe that top-activating neurons in earlier layers causally influence those in later layers. In particular, interventions on neurons in layer 5 or earlier lead to automatic changes in layer 6 activations (see Figure 5).
Figure 14: Additional activation maps of neurons that activate highly on the top outlier position. Even in layers before outliers emerge (layer 6), we observe that there exist neurons whose activation maps closely align with the output patch norms.
A.5 Full experiment results on test-time registers in OpenCLIP
@@
A.6 Additional experiments on register neurons in OpenCLIP
@@
A.7 Investigating outliers in DINOv2
@@
A.8 Extended Dense Prediction Results
@@
A.9 Additional LOST object discovery results
@@ A.10 Additional VLM results
A.11 Removing outliers with attention biases
While we focus on shifting outliers outside the image in this work, high-norm outliers still persist in the extra test-time register. Here, we propose a preliminary method to eliminate these artifacts without additional register tokens, by using test-time attention biases to fully remove high-norm outliers from the residual stream.
Incorporating attention biases at test time. Sun et al. (2024) observed that language and vision transformers have massive values at certain dimensions and proposed to remove them by incorporating attention biases during training. Specifically, given the query, key, and value matrices $Q,K,V\in\mathbb{R}^{T\times d}$ for each attention head, they train additional parameters $\mathbf{k}^{\prime},\mathbf{v}^{\prime}\in\mathbb{R}^{d}$ such that
$\text{Attention}(Q,K,V;\mathbf{k}^{\prime},\mathbf{v}^{\prime})=\text{softmax}\left(\frac{Q\begin{bmatrix}K^{T}&\mathbf{k}^{\prime}\end{bmatrix}}{\sqrt{d}}\right)\begin{bmatrix}V\\ \mathbf{v}^{\prime T}\end{bmatrix}$
We propose to create these attention biases in OpenCLIP ViT-B/16 training-free by approximating the effect of a test-time register on the attention computations. To set $\mathbf{k^{\prime}}$ and $\mathbf{v^{\prime}}$ for each attention head, we compute the mean key and value vectors of a test-time register token across 1000 images. At test-time, we zero out the activations of register neurons using the parameters from Section 3.2.
Results. We report zero-shot ImageNet performance in Table 12. Zeroing out the register neurons leads to a significant performance drop, but this drop is fully recovered by introducing the attention bias, which restores the model’s original performance. This suggests that zeroing out registers effectively removes outliers, while the attention bias approximates their influence at test time. Importantly, we no longer observe high-norm outliers across all layers (Figure 26). We also find that attention biases create clean attention maps free of artifacts comparable to using a test-time register (Figure 27). Our findings support the claim by Sun et al. (2024) that registers primarily function as attention biases. Although massive activations are relatively modest in vision transformers (e.g., < 500), they can exceed the mean activation by up to 10,000$\times$ in language models, creating challenges for quantization—where specialized methods are often needed to handle such outliers (Xiao et al. (2023)). Our results suggest that these large activations can be controlled without any additional training overhead. We leave the application of our method to the language domain for future work.
Figure 26: Averaged max patch norm across all layers in OpenCLIP-B/16 using attention biases. We find that using attention biases completely removes the presence of high-norm outliers from the residual stream.
OpenCLIP Acc.
ViT-B/16 71.3
+ zeroed register neurons 55.6
+ attention bias 71.3
Table 12: Creating attention biases at test-time maintains zero-shot ImageNet performance. Zeroing out register neurons significantly drops performance, but attention biases recover the loss.
Figure 27: Using attention biases produces similarly clean attention maps as using a test-time register. We compute the attention map of the CLS token averaged across the attention heads of the last layer on OpenCLIP-B/16, finding that both attention biases and test-time registers remove artifacts.